# Seeing through Light and Darkness: Sensor-Physics Grounded Deblurring HDR NeRF from Single-Exposure Images and Events

### 13 Mar 2026 v3

@@ 5 Experiments
Since EvHDR-NeRF [6] has not released its datasets, we construct both synthetic and real datasets for the task of HDR novel view synthesis from single-exposure blurry LDR images and corresponding events. We also adopt the public Real-World-Challenge dataset [34] with challenging blurry ERGB input and GT sharp test images to evaluate the deblurring effect of See-NeRF. Due to space limitations, we only briefly describe the collected dataset here and provide further dataset generation and experimental details in the Secs. B, C, and D of the supplementary materials.
@@ 5.1 Experiment settings
Synthetic dataset: The synthetic dataset consists of 8 HDR “Blender” scenes from HDR-NeRF [19]. Each scene contains 18 training views and 17 test views. We use Blender [1] and v2e [17] to generate blurry images with exposure time and corresponding events as in [35] for training views. We also provide sharp LDR images with exposure times for the image-based HDR NVS methods without deblurring effects. For test views, we provide GT sharp HDR images and LDR images with for HDR and novel exposure evaluation as in [19].
Real dataset: The real dataset includes 5 scenes captured under extreme lighting conditions using the DAVIS 346 event camera [41]. Each scene has 16 training views and 12 test views. For training views, we use a handheld DAVIS 346 to capture the blurry LDR images with exposure time along with the corresponding event data. We also captured sharp LDR images with a tripod with exposure times . For the test views, we use a tripod to capture GT sharp LDR images with exposure times , . We also capture the GT sharp LDR images with , , , and use the classic multi-exposure HDR reconstruction algorithm DeBevec [9] to generate the GT sharp HDR image with the 5 sharp LDR images wtih different exposures for each test view as in RawHDR [56].
@@ 5.2 HDR and novel exposure NVS experiment
We choose the image-based HDR NVS works [19, 4, 27, 49] for comparison, and use sharp single-exposure images as input, considering they cannot handle the blurry images. Additionally, we use multi-exposure images as input for HDR-NeRF as a reference. For the ERGB-based methods, we use EDI [33] and HDRev [48] to deblur the blurry LDR images and generate corresponding sharp HDR images with event data, which are input into the HDR-NeRF for training (HDR-NeRF+). We also use the latest ERGB-Based deblurring HDR NeRF methods, EvHDR-NeRF, for comparison.
Quantitative Results: As in Tab. 1, our See-NeRF significantly outperforms the compared methods on both synthetic and real datasets. The HDR reconstruction results on the real dataset are even better than HDR-NeRFref. In comparison, RGB-based HDR NVS methods exhibit constrained performance owing to their dependence on multi-exposure input. Cascaded method HDR-NeRF+ achieves superior results compared to RGB-based methods but underperforms EvHDR-NeRF due to its decoupled pipeline that separately handles 3D reconstruction and HDR sharp image reconstruction, constrained by the inherent bottlenecks of EDI and HDRev. EvHDR-NeRF realizes the second-best overall results, but is significantly worse than our See-NeRF.
Qualitative Results: Fig. 5 displays the qualitative HDR NVS results, and our See-NeRF achieves the best performance. The results of EvHDR-NeRF exhibit slight color deviation and blur artifacts, which stem from inaccurate CRF estimation and unmodeled radiometric integration of the RGB sensor. HDR-NeRF+ generates reconstructions with elevated luminance but significant color distortion, a limitation primarily attributable to HDRev’s suboptimal radiance estimation. Lack of multi-exposure input, RGB-based methods have serious color casts, especially in the overexposed parts of the real data. For the novel exposure synthesis task, See-NeRF perfectly recovers both overexposed and underexposed areas, while other methods are affected by color and brightness mismatch as in Fig. 6.
@@ 5.3 Deblurring NVS experiment
In this experiment, we compared our method with recent RGB-based and ERGB-based deblurring NVS works, and the quantitative results are shown in Tab. 2. Our See-NeRF achieves significant performance improvements, especially on the LPIPS, proving that our proposed event and RGB mapping fields effectively bridge the difference between the scene radiance and the sensor output, thereby strengthening the use of non-ideal data to reconstruct the sharp 3D scene representation. Fig. 7 shows that See-NeRF successfully recovers the blue eyes and two pink cheek dots of the rabbit, while the best ERGB-based methods E3NeRF miss these subtle details. The results of the best RGB-based method, DP-NeRF, are inferior due to the lack of event enhancement.
@@
| Event Input | Event Mapping Field | RGB Mapping Field | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | |
| See-NeRFBase | - | - | - | - | - | - | - | - | - | 28.79 | .9201 | .2939 |
@@ 5.4 Ablation study
Effect of RGB mapping field: Comparing See-NeRFBase (training the base model without events) and See-NeRFNoe (training See-NeRF without events) in Tab. 3 shows that our RGB mapping field enables the network to perform HDR NVS while improving the deblurring NVS results.
Pixel-Wise CRF versus 3D points CRF: The results of See-NeRF3D of Tab. 3 show that substituting our pixel-wise RGB mapping field with the 3D point CRF in [19, 4, 6] significantly reduces performance in HDR and deblurring NVS and leads to incorrect CRF curve estimation in Fig. 8.
Effect of event input: The second and third lines of Tab. 3 prove that event supervision plays a key role in both HDR and deblurring NVS as theoretically deduced in Sec. 3. Fig. 8 further indicates that events can help our model to learn the correct CRF curve of the RGB sensor.
Effect of event mapping field (EMF): Comparing See-NeRFNoEMF and See-NeRF in Tab. 3, our EMF marginally helps align the CRF curve (Fig. 8), achieving further improvement on HDR and Deblurring NVS. However, replacing our EMF with the eCRF model in EvDeblur-NeRF [5] leads to barely satisfactory reconstruction results and CRF estimation (See-NeRFEvD in Tab. 3 and Fig. 8).
| Robustness Different Input | Robustness to Event Noise | ||||||
| Single-Exp | PSNR↑ | SSIM↑ | LPIPS↓ | Noise Level | PSNR↑ | SSIM↑ | LPIPS↓ |
@@ 5.5 Discussions
Robustness to LDR images with different exposures and event noise: As shown in Tab. 4, we conduct a robustness experiment on the synthetic “Catroom” scene for the HDR NVS task. Firstly, we use the single-exposure images with as inputs, respectively. The results show that single-exposure inputs with longer exposure times yield better results, indicating that the low-light regions of the scene are more difficult to learn than the high-light regions. Secondly, we use events with different noise levels as input. Tab. 4 and the supplement video results on real data show that event noise does not lead to a significant performance degradation, because NeRF training refers to multi-view input, which can inherently weaken the impact of sparse random noise in each individual view, as proved in RawNeRF [30]. Therefore, See-NeRF does not target the processing of event noise, which could be a future optimization direction of ERGB-based NVS methods.
| Novel Exposure | HDR | Time | |||||
| PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | hours | |
| HDR-GS | 19.19 | .7745 | .2537 | 9.25 | .4036 | .6376 | 0.03 |
| Ours + 3DGS | 20.39 | .8204 | .2166 | 18.27 | .7244 | .3481 | 0.05 |
| Ours | 24.15 | .8576 | .1154 | 19.43 | .7979 | .2469 | 2.5 |
Training time and availability on 3DGS: The training time of our See-NeRF (2.5 h) is slightly longer than that of HDR-NeRF (2.1 h) due to the introduction of event-related computation. We also replace the NeRF network of our method with 3DGS and conduct an experiment on the “computerhard” scene of our real data. As in Tab. 5, “Ours + 3DGS” is better than HDR-GS, especially on the HDR NVS task, and the training speed is also increased by about 50 times with the high-speed rendering of 3DGS. However, the performance of “Ours + 3DGS” is significantly lower than See-NeRF on the novel exposure task. We think this is because the MPL-based camera mapping fields do not perfectly match the 3DGS during optimization, and ERGB-based HDR 3DGS may be a promising future work.
@@ 6 Conclusion
In this paper, we propose a unified Sensor-Physics grounded framework, See-NeRF, for deblurring HDR NVS from single-exposure LDR images and corresponding events, which bridges the scene radiance and the sensor-recorded ERGB data based on the physical imaging laws and achieves the state-of-the-art deblurring HDR NVS results.
A Details of events calibration and division
Events calibration function : For the input event bin between and , we use and to represent the first and last triggered event at pixel respectively. Then the offset of event set can be expressed as:
where represents the function that if we can not find the or (there is no event triggered in the previous event bin or the next event bin at pixel ), the output is . Since this operation is the same on all pixels, we omit the in Eq. 16.
Events division: For each training view, we split the input event stream that corresponding to the input LDR blurry image into bins , where and are the start and end exposure time of the image. are the time points that divide the event stream into event bins with an equal number of events as in the E3NeRF [34], which leverages the temporal blur prior in events distribution into event loss optimization to achieve better results.
@@ B.1 Synthetic dataset:
Our synthetic dataset consists of 8 HDR Blender [1] scenes (“bathroom”, “catroom”, “dogroom”, “diningroom”, “sofa”, “sponza”, “toyroom”, “warmroom”) in the synthetic data of HDR-NeRF [19], which have both highlight and darkness areas in each scene, resulting in the captured single-exposure images containing both overexposure and underexposure areas as in the first column of Fig. 12. All the data are at a resolution of . Below, we introduce the training and test data generation process, respectively.
Training images and events generation: For each scene, there are 18 views of LDR blurry images with a single exposure time and corresponding events within the exposure time for training. For each view, we add camera shaking and generate 17 sharp HDR raw images at time with camera poses . The camera pose changes over time, indicating the camera shake process. Note that , where and are the start time and end time of the exposure as defined in the main paper. Then the 17 sharp HDR raw images are averaged to obtain the blurred HDR raw image . We used the classic Reinhard tonemapping [38] algorithm in Eq. 17 to convert the blurred HDR raw image into an LDR single-exposure blurred image with exposure time for training, which is the same as in HDR-NeRF [19]:
@@
Test images generation: For each scene, there are 17 novel views for testing. For each test view, we render 1 ground-truth (GT) sharp HDR raw image with Blender for the HDR novel view synthesis (NVS) test. We also use Eq. 17 to transform the sharp HDR image into two GT sharp LDR images at novel exposure times and for the novel exposure NVS test.
@@ B.2 Real dataset:
Our synthetic dataset consists of 5 real-world extreme lighting scenes (“bear”, “computer”, “computerhard”, “desktop”, “table”) with both highlight and darkness areas, resulting in the captured single-exposure LDR images containing both overexposure and underexposure areas. The upper part of Tab. 6 shows the minimum and maximum luminance of the scenes in our real data, which contains a very wide dynamic range. All data is obtained by a DAVIS 346 Color camera [41], which can capture spatial-temporal aligned images and events at resolution .
Training images and events generation: For each scene, there are 16 training views. For each view, we capture 1 LDR blurry image at a single exposure time and corresponding events within the exposure time with a handheld DAVIS 346 Color camera.
Additional training images: For each training view, we also capture three sharp LDR images at exposure times , , and with a tripod-fixed DAIVS 346 Color camera to train the RGB-based HDR NVS works [19, 4, 27, 49] without deblurring effects and to train HDR-NeRF [19] with multi-exposure sharp LDR images as input for reference.
Test images generation: For each scene, there are 12 novel views for testing. For each test view, we capture 5 LDR sharp images at exposure times with a tripod-fixed DAIVS 346 Color camera. The LDR sharp images with novel exposure times and are used for the novel exposure NVS test, which is the same as the synthetic dataset. The GT HDR sharp image for the HDR NVS test is obtained by merging the five LDR sharp images with exposure times with the classic multi-exposure HDR reconstruction algorithm Debevec [9] as in RawHDR[56].
@@ C Implementation details
Our framework is built upon E2NeRF [35], and all experiments are conducted on a single RTX 3090 GPU. We set the number of sampled points on each ray to 64 and 128 for the coarse and fine networks, respectively. The batch size of sampled rays is set to 512 and 1024 for the real data and synthetic data, respectively, and we take 50k iterations for each scene. The above NeRF network settings are the same for all the compared NeRF-based methods and our See-NeRF in the experiments. For the 3DGS-based methods, the number of iterations is set to 30000, and other parameters are set to their default parameters for each method. For the detailed evaluation of hyperparameters , , and training time of our See-NeRF, please refer to Sec. D.2.

### 02 Apr 2026 v4

@@ 5 Experiments
Since EvHDR-NeRF [6] has not released its datasets, we construct both synthetic and real datasets for the deblurring HDR NVS task from single-exposure blurry LDR images and corresponding events. We also adopt the public Real-World-Challenge dataset [36] with challenging blurry ERGB input and GT sharp test images to evaluate the deblurring effect of See-NeRF. Due to space limitations, we only briefly describe the collected dataset here and provide further dataset generation and experimental details in the Secs. B, C, and D of the supplementary materials.
@@ 5.1 Experiment settings
Synthetic dataset: The synthetic dataset consists of 8 HDR “Blender” scenes from HDR-NeRF [19]. Each scene contains 18 training and 17 test views. We use Blender [1] and v2e [17] to generate blurry images with exposure time and corresponding events as in [37] for training views. We also provide sharp LDR images with exposure times for the RGB-based HDR NVS methods without deblurring effects. For test views, we provide GT sharp HDR images and LDR images with for HDR and novel exposure evaluation as in [19].
Real dataset: The real dataset includes 5 scenes captured under extreme lighting conditions using the DAVIS 346 event camera [45]. Each scene has 16 training and 12 test views. For training views, we use a handheld DAVIS 346 to capture the blurry LDR images with exposure time along with the corresponding event data. We also captured sharp LDR images with a tripod with exposure times . For the test views, we use a tripod to capture GT sharp LDR images with exposure times , . We also capture the GT sharp LDR images with , , , and use the classic multi-exposure HDR reconstruction algorithm DeBevec [9] to generate the GT sharp HDR image with the 5 sharp LDR images with different exposures for each test view as in RawHDR [60].
@@ 5.2 HDR and novel exposure NVS experiment
We choose the RGB-based HDR NVS works [19, 4, 29, 53] for comparison with sharp single-exposure images as input, considering they cannot handle the blurry images. We also use multi-exposure images as input for HDR-NeRF as a reference (HDR-NeRFref). For the ERGB-based methods, we use EDI [35] and HDRev [52] to deblur the blurry LDR images and generate sharp HDR images with event data, which are input into the HDR-NeRF for training (HDR-NeRF+). We also compare with the ERGB-based deblurring HDR NeRF method, EvHDR-NeRF.
Quantitative results: As in Tab. 1, our See-NeRF significantly outperforms the compared methods on both synthetic and real datasets. The HDR reconstruction results on the real dataset are even better than HDR-NeRFref. In comparison, RGB-based HDR NVS methods exhibit constrained performance owing to their dependence on multi-exposure input. Cascaded method HDR-NeRF+ achieves superior results compared to RGB-based methods but underperforms EvHDR-NeRF due to its decoupled pipeline that separately handles 3D reconstruction and HDR sharp image generation, constrained by the inherent bottlenecks of EDI and HDRev. EvHDR-NeRF realizes the second-best overall results, but is significantly worse than our See-NeRF.
Qualitative results: Fig. 5 displays the qualitative HDR NVS results, and our See-NeRF achieves the best performance. The results of EvHDR-NeRF exhibit slight color deviation and blur artifacts, which stem from inaccurate CRF estimation and unmodeled radiance integration of the RGB sensor. HDR-NeRF+ generates results with elevated luminance but significant color distortion, a limitation primarily attributable to suboptimal radiance estimation of HDRev. Lack of multi-exposure input, RGB-based methods have serious color casts, especially in the overexposed parts of the real data. For the novel exposure NVS task, See-NeRF perfectly recovers both overexposed and underexposed areas, while other methods are affected by color and brightness mismatch as shown in Fig. 6.
@@ 5.3 Deblurring NVS experiment
In this experiment, we compared our method with recent RGB-based and ERGB-based deblurring NVS works, and the quantitative results are shown in Tab. 2. Our See-NeRF achieves significant performance improvements, especially on the LPIPS, proving that our proposed event and RGB CRF models effectively bridge the sensor-physics discrepancy, strengthening the use of non-ideal data to reconstruct the sharp 3D scene. Fig. 7 shows that See-NeRF successfully recovers the blue eyes and two pink cheek dots of the rabbit, while the best ERGB-based deblurring NVS method E3NeRF misses these subtle details. The results of the best RGB-based deblurring NVS method, DP-NeRF, are inferior due to the lack of event enhancement.
@@
| Event Input | Event CRF Model | RGB CRF model | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | |
| See-NeRFBase | - | - | Naive [33, 37, 36] | - | - | - | - | - | - | 28.79 | .9201 | .2939 |
@@ 5.4 Ablation study
Effect of RGB CRF model: Comparing See-NeRFBase (training the base model without events) and See-NeRFNoEv (training See-NeRF without events) in Tab. 3 shows that our RGB CRF model enables the network to perform HDR NVS while improving the deblurring NVS results.
2D pixel-wise CRF versus 3D points CRF: The results of See-NeRF3D of Tab. 3 show that substituting our 2D pixel-wise RGB CRF model with the 3D point CRF of [19, 4, 6] significantly reduces performance in HDR and deblurring NVS and leads to incorrect CRF curve estimation in Fig. 8.
Effect of event input: The second and third lines of Tab. 3 prove that event supervision plays a key role in both HDR and deblurring NVS, as theoretically deduced in Sec. 3. Fig. 8 further indicates that events can help our model to learn the correct CRF curve of the RGB sensor.
Effect of event CRF model: Comparing See-NeRFNoEM and See-NeRF in Tab. 3, our event CRF model marginally helps align the CRF curve (Fig. 8), achieving further improvement on HDR and Deblurring NVS. However, replacing our event CRF model with the eCRF in EvDeblur-NeRF [5] leads to barely satisfactory reconstruction results and CRF estimation (See-NeRFEvD in Tab. 3 and Fig. 8).
| Robustness to Different Input | Robustness to Event Noise | ||||||
| Single-Exposure | PSNR↑ | SSIM↑ | LPIPS↓ | Noise Level | PSNR↑ | SSIM↑ | LPIPS↓ |
@@ 5.5 Discussions
Robustness to LDR images with different exposures and event noise: As shown in Tab. 4, we conduct a robustness experiment on the synthetic “Catroom” scene for the HDR NVS task. Firstly, we use the single-exposure images with exposure times as inputs, respectively. The overall results do not fluctuate significantly, while longer exposure input leads to better results, indicating that the low-light regions of the scene are more difficult to learn than the highlight regions. Secondly, we use events with different noise levels as input. Tab. 4 and the supplement video results on real data show that event noise does not lead to a significant performance degradation, because NeRF training refers to multi-view input, which can inherently weaken the impact of sparse random noise in each individual view, as proved in RawNeRF [32]. Therefore, See-NeRF does not target the processing of event noise, and it could be a future optimization direction of ERGB-based NVS methods.
| Novel Exposure | HDR | Training | Rendering | |||||
| PSNR↑ | SSIM↑ | LPIPS↓ | PSNR↑ | SSIM↑ | LPIPS↓ | hours | fps | |
| HDR-GS [4] | 19.19 | .7745 | .2537 | 9.25 | .4036 | .6376 | 0.03 | 4.29 |
| Ours + 3DGS | 20.39 | .8204 | .2166 | 18.27 | .7244 | .3481 | 0.05 | 5.11 |
| Ours | 24.15 | .8576 | .1154 | 19.43 | .7979 | .2469 | 2.5 | 0.37 |
Training and rendering time and availability on 3DGS: The training time of our See-NeRF (2.5 h) is slightly longer than that of HDR-NeRF (2.1 h) due to the introduction of event-related computation. We also replace the NeRF network of our method with 3DGS and conduct an experiment on the “computerhard” scene of our real data. As shown in Tab. 5, “Ours + 3DGS” is better than HDR-GS, especially on the HDR NVS task, and the training speed is also increased by about 50 times with the high-speed rendering of 3DGS. Note that the rendering speed of “Ours + 3DGS” is even faster than HDR-GS, demonstrating our 2D pixel-wise RGB CRF model is more efficient than the 3D point CRF of HDR-GS. However, the performance of “Ours + 3DGS” is significantly lower than See-NeRF on the novel exposure task. We think this is because the MLP-based CRF model does not perfectly match the 3DGS during optimization, and ERGB-based HDR 3DGS may be a promising future work.
@@ 6 Conclusion
In this paper, we propose a unified sensor-physics grounded framework, See-NeRF, for deblurring HDR NVS from single-exposure LDR images and corresponding events, which bridges the discrepancy between the scene radiance and the sensor-recorded ERGB data based on the physical imaging laws and the sensor processing characteristics, achieving the state-of-the-art deblurring HDR NVS results.
Acknowledgements
This work is partially supported by grants from the National Natural Science Foundation of China (No. 62132002), Guizhou Provincial Major Scientific and Technological Program (Qiankehe Zhongda [2025] No. 032), Beijing Nova Program (No.20250484786), the Fundamental Research Funds for the Central Universities, and the China Scholarship Council (No. 202506020106)
A Details of event calibration and division
Photometric quantity calibration function : For the input event bin between and , we use and to represent the first and last triggered event at pixel , respectively. Then we can calculate the offset of each pixel for the input event bin with:
where is a conditional function to determine whether there are adjacent events triggered before and after the current event bin, and whether the polarities are consistent:
Since the above operations are the same for all pixels, we omit the in Eq. 16 and Eq. 17.
Events division: For each training view, we split the input event stream that corresponding to the input LDR blurry image into bins , where and are the start and end exposure time of the image. are the time points that divide the event stream into event bins with an equal number of events as in the E3NeRF [36], which leverages the temporal blur prior in event distribution for event loss optimization, achieving better performance.
@@ B.1 Synthetic dataset:
Our synthetic dataset consists of 8 HDR Blender [1] scenes (“bathroom”, “catroom”, “dogroom”, “diningroom”, “sofa”, “sponza”, “toyroom”, “warmroom”) in the synthetic data of HDR-NeRF [19], which contain both highlight and darkness areas in each scene, resulting in the rendered single-exposure images containing both overexposure and underexposure areas as in the first column of Fig. 12. All synthetic data are at a resolution of . Below, we introduce the training and test data generation process, respectively.
Training images and events generation: For each scene, there are 18 views of LDR blurry images with a single-exposure time and corresponding events within the exposure time for training. For each view, we add camera shaking and generate 17 sharp HDR raw images at time with camera poses . The camera pose changes over time, indicating the camera shake process. Note that , where and are the start time and end time of the exposure as defined in the main paper. Then the 17 sharp HDR raw images are averaged to obtain the blurred HDR raw image . We used the classic Reinhard tone mapping [40] algorithm in Eq. 18 to convert the blurred HDR raw image into an LDR blurred image with exposure time for training as in HDR-NeRF [19]:
@@
Test images generation: For each scene, there are 17 novel views for testing. For each test view, we render 1 ground-truth (GT) sharp HDR raw image with Blender for the HDR novel view synthesis (NVS) test. We also use Eq. 18 to transform the sharp HDR image into two GT sharp LDR images with novel exposure times and for the novel exposure NVS test.
@@ B.2 Real dataset:
Our real dataset consists of 5 real-world extreme lighting scenes (“bear”, “computer”, “computerhard”, “desktop”, “table”) with both highlight and darkness areas, resulting in the captured single-exposure LDR images containing both overexposure and underexposure areas. The upper part of Tab. 6 shows the minimum and maximum luminance of the scenes in our real data, which contains a very wide dynamic range. All data is obtained by a DAVIS 346 Color camera [45], which can capture spatial-temporal aligned images and corresponding events at resolution .
Training images and events generation: For each scene, there are 16 training views. For each view, we capture 1 LDR blurry image at a single-exposure time and corresponding events within the exposure time with a handheld DAVIS 346 Color camera.
Additional training images: For each training view, we also capture three sharp LDR images with exposure times , , and with a tripod-fixed DAIVS 346 Color camera to train the RGB-based HDR NVS works [19, 4, 29, 53] without deblurring effects and to train HDR-NeRF [19] with multi-exposure sharp LDR images as input for reference.
Test images generation: For each scene, there are 12 novel views for testing. For each test view, we capture 5 LDR sharp images with exposure times with a tripod-fixed DAIVS 346 Color camera. The LDR sharp images with novel exposure times and are used for the novel exposure NVS test, which is the same as the synthetic dataset. The GT HDR sharp image for the HDR NVS test is obtained by merging the five LDR sharp images with exposure times with the classic multi-exposure HDR reconstruction algorithm Debevec [9] as in RawHDR [60].
@@ C Implementation details
Our framework is built upon E2NeRF [37], and all experiments are conducted on a single RTX 3090 GPU. We set the number of sampled points on each ray to 64 and 128 for the coarse and fine networks, respectively. The batch size of sampled rays is set to 512 and 1024 for the real data and synthetic data, respectively, and we take 50k iterations for each scene. The above NeRF network settings are the same for all the compared NeRF-based methods and our See-NeRF in the experiments. For the 3DGS-based methods, the number of iterations is set to 30000, and other parameters are set to their default parameter. For the detailed evaluation of hyperparameters , , and training time of our See-NeRF, please refer to Sec. D.2.